# RAISE: Enhancing Scientific Reasoning in LLMs via Step-by-Step Retrieval

### v1

@@ Abstract
Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings. To deal with these challenges for scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from in-the-wild corpus. RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval. We observe that RAISE consistently outperforms other baselines on scientific reasoning benchmarks. We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant.
@@
Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University {minhaeoh, junglee,lnktop}@snu.ac.kr, jeonghye.kim@kaist.ac.kr, {donggunseo,kimtaeuk}@hanyang.ac.kr
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities for diverse scientific domains that require intricate reasoning and the integration of diverse knowledge sources (Zhang et al., 2024c, a; Prabhakar et al., 2025; Rueda et al., 2025). In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge. To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation. For step-wise reasoning, it has been shown that solving reasoning problems through step-wise reasoning process is more effective than attempting to solve them in a single step, similar to how humans approach challenging problems Wei et al. (2022); Zhao et al. (2024a); Jin et al. (2024); Lai et al. (2024); Yao et al. (2023). For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucination and deal with timely knowledge, which is especially important for science domain Lewis et al. (2020b); Bang et al. (2025); Zhong et al. (2025); Xiong et al. (2024).
To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning on various domains (Jeong et al., 2024; Zhang et al., 2025a). However, most of the works focus on relatively simple multi-hop question answering tasks or non-scientific domains (Guan et al., 2025; Jin et al., 2025; Wang et al., 2025) or assume task-related retrieval corpus which is hard to obtain compared to in-the-wild retrieval source such as Wikipedia (Zhong et al., 2025; Qian et al., 2023; Anand et al., 2023). Solving challenging scientific reasoning tasks such as graduate-level biology, physics, etc, using an in-the-wild corpus is very challenging since merely retrieving superficial knowledge is insufficient for these tasks. Instead, the retrieved information should contain relevant logics needed to solve the problem (Rueda et al., 2025). Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem. Without considering the evolving information needed for each reasoning process, RAG might even deteriorate the downstream task performance. The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
@@
To address these challenges, we introduce RAISE (step-by-step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning. Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval. For problem decomposition, LLMs decompose the original question into subquestions along with search queries that can assist in solving each reasoning step. Then, for each search query, LLMs convert the search query into a logical query, which is a query that contains logical explanations. Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrievers to retrieve based on logic similarity rather than superficial domain similarity. Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step. Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios.
We evaluate our framework on three challenging scientific reasoning benchmarks: GPQA Rein et al. (2024), SuperGPQA Team et al. (2025), and MMLU Hendrycks et al. (2021); Wang et al. (2024b).RAISE consistently outperforms various baselines approaches that include applying RAG or problem decomposition. This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step. Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity. Our code is available at https://github.com/minhaeoh/RAISE.
@@ 2 Preliminary
Step-by-Step Reasoning in LLMs.    LLMs are capable of performing multi-step reasoning over complex input queries by internally chaining intermediate inferences. This step-by-step reasoning process involves decomposing a question into sub-problems, maintaining coherence across steps, and generating a final answer. Formally, given a query $x$, the model implicitly constructs a latent reasoning trajectory $\{r_{t}\}_{t=1}^{T}$, and generates the answer $y$ conditioned on this chain:
@@
RAG for Single-Step Reasoning.    We address the task of generating a response $y$ given an input $x$, enhanced by retrieval from an external corpus $\mathcal{D}$. RAG combines a retriever and a generator to condition the output on both the input and relevant documents.
@@
Top-$k$ documents are retrieved, and a generator (e.g., BART Lewis et al. (2020a), T5 Raffel et al. (2020)) produces $y$ based on both $x$ and $d_{i}$.
@@ 3 RAISE
We propose RAISE (step-by-step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval. Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each. The overview of RAISE is provided in Figure 1 and Algorithm 1. The method consists of three main stages:
@@
Problem Decomposition.    RAISE begins with the decomposition of a scientific reasoning problem $x$ into a sequence of intermediate subquestions $\{r_{i}\}_{i=1}^{n}$, each corresponding to a distinct reasoning step. This planning process is performed by prompting a language model $P_{\theta}$ with $x$ and prompt $p_{1}$ to identify the logical structure required to solve the problem. Formally, we generate:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1})$ (1)
@@
Logical Query Generation.    In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$. Search queries $q_{i}$ are short key phrases, where the logical process is absent. Therefore, using $q_{i}$ for retrieval will focus on domain similarity instead of logical similarity. Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion. The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g. add essential scientific or mathematical explanation the search result would contain):
@@
Using only the subquestion $r_{i}$ to generate a query often results in noisy or overly specific formulations that include problem-specific variables or conditions irrelevant to general knowledge retrieval. On the other hand, using only the initial query $q_{i}$ lacks sufficient context to capture the intended reasoning path. By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$. Even if the reformulated query $q_{i}^{*}$ contains facutal inaccuracies, it tends to retrieve passages from corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem. Figure 2 presents example queries generated by RAISE ,Step-back and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aglined with the reasoning intent.
Logical Retrieval.    The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
@@
Here, $\mathcal{R}$ is a top-$k$ retrieval function over in-the-wild corpus $\mathcal{C}$ (e.g., Wikipedia), and $D_{i}=\{d_{i1},\dots,d_{ik}\}$ denotes the retrieved documents. The goal is not to retrieve direct answers but to obtain passages that contain relevant background knowledge or reasoning procedures necessary to solve $r_{i}$. Also, since irrelevant documents might be also retrieved, we set a retrieval threshold $T$, where we retrieve documents that have higher retrieval similarity than $T$.
@@
Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
Direct Reasoning Decomposed Reasoning
CoT PS Step-back CoT+RAG Least-to-Most Step-back RAG RAISE
GPQA
Overall 42.42 39.39 43.94 45.96 44.95 44.44 45.95 51.01
SuperGPQA
science-hard 7.54 9.05 9.05 7.54 6.03 5.03 6.03 10.05
science-middle 15.08 14.07 17.59 12.56 14.57 15.08 15.08 19.60
engineering-hard 7.54 4.02 5.03 7.54 10.05 7.54 8.04 10.55
MMLU
(Pro) Chemistry 25.44 26.50 26.86 25.18 24.56 22.70 22.97 28.36
(STEM) College Chemistry 49.50 45.00 44.00 43.00 45.40 43.00 46.00 51.00
@@ 4.1 Experimental Setup
Datasets.    To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning. We provide more details on the evaluation dataset in Appendix A.
@@
  • SuperGPQA (Team et al., 2025): This dataset evaluates the knowledge and reasoning abilities of Large Language Models across 285 graduate-level disciplines. We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
  • MMLU (Hendrycks et al., 2021; Wang et al., 2024b): Among various domains and difficulty levels of MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024b), we focus on chemistry domain since it requires step-by-step reasoning with knowledge of specialized terminology.
Retriever Configuration.    We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
Language Model.    For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks. For the remaining tasks, we use the smaller LLaMA 3.1-8B model Grattafiori et al. (2024) due to computational constraints, applying it to SuperGPQA, and MMLU. We provide detailed hyperparameters and experiment setups in Appendix B.
4.2 Baselines
To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with two groups of baselines: Direct Reasoning and Decomposed Reasoning. Direct Reasoning methods solve the problem in a single inference step without decomposing it into subquestions. This group includes three prompt-based approaches and one retrieval-augmented method. The three prompting methods-Chain-of-Thought (CoT) Wei et al. (2022), Plan-and-Solve (PS) Wang et al. (2023), Step-back Zheng et al. (2024) -differ in how the prompt guides the language models’ reasoning. Additionally, we include CoT+RAG Lewis et al. (2020b), a conventional retrieval-augmented method that retrieves documents using the original question as query.
Decomposed Reasoning methods first break down the original problem into subquestions and solve them sequentially. This group includes Least-to-Most Zhou et al. (2023), which solves each subquestion directly without retreival, Step-back, which applies principle extraction at the subquestion level, and RAG, which retrieves and uses evidence for each subquestion individually. More detailed descriptions of each baseline can be found in Appendix B.1
@@ 4.3 Main Results
As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets: GPQA, SuperGPQA, and MMLU. On GPQA, RAISE achieves 51.01%, surpassing the next-best method (CoT+RAG) by over 5 percentage points. On SuperGPQA, which consists of especially challenging science and engineering questions, RAISE achieves the highest performance across all subdomains, with gains especially notable in the science-middle subset (19.60%, +2 pts over the best baseline). On MMLU, RAISE also shows significant improvement, particularly in (Pro) Chemistry (28.36%) and STEM College Chemistry (51.00%), indicating its robustness across both professional and academic scientific reasoning tasks. This result shows that RAISE effectively decomposes scientific reasoning questions and retrieves documents that help the model to solve the questions.
@@
Unlike RAISE, decomposed reasoning methods do not always lead to better performance. This is often due to a lack of logically relevant knowledge especially for smaller LLMs, which limits their ability to reason effectively across multiple steps (Hosseini et al., 2024; Xu et al., 2025). In fact, the original papers of our baseline models that use decomposed reasoning mostly focused on proprietary large LLMs while we focus on open-sourced smaller LLMs. Despite well-structured decomposition into subquestions, a lack of logically relevant knowledge may cause the model to produce hallucinations or unsupported reasoning steps. Especially for smaller models that lack sufficient prior knowledge, problem decomposition can become counterproductive unless logically relevant documents are retrieved to support the reasoning process.
@@
Also, we observe that in some cases, models without retrieval outperform their retrieval-augmented counterparts. While retrieval is generally expected to enhance performance by providing additional information, these results reveal that naively applying retrieval can be ineffective or even harmful. This issue becomes particularly pronounced in settings involving open-domain or in-the-wild corpora, where the risk of retrieving noisy or irrelevant content is high. Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning. These findings highlight the importance of retrieving logically relevant knowledge, and further underscore the effectiveness of our method.
Table 2: Comparison of decomposed reasoning strategies with retrieval. RAISE consistently outperforms baselines across datasets.
GPQA SuperGPQA MMLU
RAG 45.95 9.72 24.67
Step-back+RAG 43.43 10.05 33.03
HyDE 28.28 9.22 35.99
RAISE 51.01 13.57 39.68
4.4 Additional Results
Impact of Logical Query Generation.    To evaluate the impact of logical query generation in retrieval-augmented reasoning, we conduct an experiment comparing our method, RAISE, with three RAG-based decomposed reasoning baselines: RAG, step-back+RAG (Zheng et al., 2024), and HyDE (Gao et al., 2023a). While all methods incorporate problem decomposition and retrieval, they differ in how the retrieval query is constructed.
RAISE begins with a short key phrase known as the initial query, and expands it into a logical query that captures not only domain similarity but also the underlying reasoning intent of the subquestion. In contrast, the standard RAG baseline directly uses the subquestion itself as the retrieval query. The step-back+RAG approach attempts to abstract a general principle from the subquestion and uses that principle as the query. HyDE generates a synthetic passage in response to the subquestion and uses this generated text as the retrieval query, aiming to enrich the query context.
As shown in Table 2, RAISE outperforms all other RAG-based baselines across all benchmarks. These results confirm that RAISE’s logical query generation goes beyond retrieving documents that are merely domain-relevant or superficially similar, enabling the retrieval of knowledge that is logically aligned with the reasoning required to solve the problem. We conduct further analysis in Section 4.5. For SuperGPQA and MMLU, Table 2 reports average scores across subsets; full per-subset results are provided in the Appendix 5.
Table 3: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1. RAISE consistently shows the best performance across all settings.
@@ 4.4 Additional Results
Step-back (Decomp.) 28.28 42.42 44.44
RAG 24.24 42.93 45.95
RAISE 30.30 47.98 51.01
Impact of Problem Decomposition.    To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3. This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE. These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway.
Applying RAISE to various LLMs.    To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B (Grattafiori et al., 2024) and GPT-4o mini (OpenAI, 2024), in addition to Mistral(used in our main experiments). As shown in Table 3, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B. This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
@@
How does the scalar potential (V) change when considering the retarded time $t_{r}$? Compensating variation in economics … additional money needed after a price change. The calculation of the retarded time $t_{r}=t-|\vec{r}-\vec{r^{\prime}}|/c$ is a speed-distance-time relation. The field observed at time $t$ must have been emitted at this earlier time. The RAISE-retrieved document directly explains the concept of retarded time in electrodynamics as the time delay due to the finite propagation speed of electromagnetic fields, which is essential for deriving the scalar potential $V(\vec{r},t)$ of a moving charge.

### v2

@@ Abstract
Scientific reasoning requires not only long-chain reasoning processes, but also knowledge of domain-specific terminologies and adaptation to updated findings. To address these challenges in scientific reasoning, we introduce RAISE, a step-by-step retrieval-augmented framework which retrieves logically relevant documents from an in-the-wild corpus. RAISE is divided into three steps: problem decomposition, logical query generation, and logical retrieval. We observe that RAISE consistently outperforms other baselines on scientific reasoning benchmarks. We analyze that unlike other baselines, RAISE retrieves documents that are not only similar in terms of the domain knowledge, but also documents logically more relevant.
@@
Minhae Oh1, Jeonghye Kim2, Nakyung Lee1, Donggeon Seo3, Taeuk Kim3, Jungwoo Lee1, 1Seoul National University, 2KAIST, 3Hanyang University
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities in scientific domains that require intricate reasoning and the integration of diverse knowledge sources Zhang et al. (2024c, a); Prabhakar et al. (2025); Rueda et al. (2025). In contrast to many other domains, scientific reasoning is especially demanding due to the need for advanced reasoning skills, knowledge of specialized terminology, and awareness of continually evolving knowledge.
To enhance the capability of LLMs to solve scientific reasoning problems, two approaches are widely used: (1) step-wise reasoning, and (2) retrieval-augmented generation. For step-wise reasoning, previous studies Wei et al. (2022); Zhao et al. (2024a); Jin et al. (2024); Lai et al. (2024); Yao et al. (2023) have shown that solving complex problems through a step-by-step process is more effective than attempting to solve them in a single step, mirroring how humans approach challenging problems. For retrieval-augmented generation (RAG), it is known that retrieving relevant external knowledge can mitigate hallucinations and help models adapt to newly emerging information, which is especially important in the scientific domain Lewis et al. (2020b); Bang et al. (2025); Zhong et al. (2025); Xiong et al. (2024).
To utilize the benefits of both worlds, recent works have adopted RAG for step-wise reasoning across various applications Jeong et al. (2024); Zhang et al. (2025a). However, most studies focus on relatively simple multi-hop question answering tasks or non-scientific domains Guan et al. (2025); Jin et al. (2025); Wang et al. (2025), or assume the availability of task-specific retrieval corpora, which are more difficult to obtain than widely accessible web-based sources such as Wikipedia Zhong et al. (2025); Qian et al. (2023); Anand et al. (2023). Solving challenging scientific reasoning tasks,such as graduate-level biology or chemistry, using an in-the-wild corpus is difficult since merely retrieving superficial knowledge is insufficient. Instead, the retrieved information should contain relevant logical connections needed to solve the problem Rueda et al. (2025). Moreover, the knowledge required for each intermediate step can vary significantly even within the same problem. Without considering the evolving information needed for each reasoning process, RAG might even deteriorate the downstream task performance. The question of what to search for and how to retrieve the appropriate external knowledge for each step when solving scientific reasoning tasks is underexplored.
@@
To address these challenges, we introduce RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented framework tailored for step-wise scientific reasoning. Our framework is divided into three steps: (1) problem decomposition, (2) logical query generation, and (3) logical retrieval. For problem decomposition, LLMs break down the original question into subquestions along with search queries that can assist in solving each reasoning step. Then, for each search query, LLMs transform it into a logical query—one that includes the underlying logical explanations. Since logical queries contain more information about what external knowledge is needed to solve the problem, they enable retrieval based on logical rather than superficial domain similarity. Finally, we perform retrieval separately for each step, obtaining documents that are logically relevant to the particular step. Instead of assuming task-relevant or well-curated retrieval source, such as question-answer pool of relevant domains, we retrieve from in-the-wild source such as Wikipedia, which enables applying to challenging real-world scenarios.
We evaluate our framework on three challenging scientific reasoning benchmarks: GPQA Rein et al. (2024), SuperGPQA Team et al. (2025), and MMLU Hendrycks et al. (2021); Wang et al. (2024b). Across these datasets, RAISE results in consistent performance improvement compared to various baseline approaches that include applying RAG or problem decomposition. This indicates that our framework effectively retrieves relevant documents that are needed to solve scientific reasoning tasks step-by-step. Through further analysis, we observe that RAISE retrieves more logically relevant documents for each step compared to baselines, highlighting that logical similarity is crucial for these challenging reasoning tasks in addition to domain similarity.
@@ 2 Preliminary
Step-by-Step Reasoning in LLMs.
LLMs are capable of performing multi-step reasoning over complex input queries by internally chaining intermediate inferences. This step-by-step reasoning process involves decomposing a question into sub-problems, maintaining coherence across steps, and generating a final answer. Formally, given a query $x$, the model implicitly constructs a latent reasoning trajectory $\{r_{t}\}_{t=1}^{T}$, and generates the answer $y$ conditioned on this chain:
@@
RAG for Single-Step Reasoning.
We address the task of generating a response $y$ given an input $x$, enhanced by retrieval from an external corpus $\mathcal{D}$. RAG combines a retriever and a generator to condition the output on both the input and relevant documents.
@@
[…]
Retrieval in In-the-Wild Settings.
We use the term “in-the-wild” to refer to open-domain corpora like Wikipedia that are not tailored for specific tasks or domains. Unlike curated corpora, they require retrieving logically relevant evidence from a large, diverse, and often tangential pool of content, making retrieval and reasoning more challenging.
@@ 3 RAISE
We propose RAISE (Step-by-Step Retrieval-Augmented Inference for Scientific rEasoning), a retrieval-augmented generation framework for scientific reasoning designed to support multi-step reasoning through fine-grained, step-aware retrieval. Rather than retrieving documents based on the entire problem as a single query, RAISE decomposes the problem into intermediate reasoning steps and performs targeted retrieval for each. The overview of RAISE is provided in Figure 1 and Algorithm 1 in Appendix A. The method consists of three main stages:
@@
Problem Decomposition.
RAISE begins with the decomposition of a scientific reasoning problem $x$ into a sequence of intermediate subquestions $\{r_{i}\}_{i=1}^{n}$, each corresponding to a distinct reasoning step. This planning process is performed by prompting a language model $P_{\theta}$ with $x$ and prompt $p_{1}$ to identify the logical structure required to solve the problem. Formally, we generate:
$\{(r_{i},q_{i})\}_{i=1}^{n}\sim P_{\theta}(\cdot\mid x,p_{1}).$ (1)
@@
Logical Query Generation.
In the second stage, each initial search query $q_{i}$ and its corresponding subquestion $r_{i}$ are jointly used to generate a logically enriched logical query $q_{i}^{*}$. Search queries $q_{i}$ are short key phrases, where the logical process is absent. Therefore, using $q_{i}$ for retrieval will focus on domain similarity instead of logical similarity. Instead, this step constructs a logically coherent query that aligns with the reasoning intent of the subquestion. The model is prompted with both $q_{i}$ and $r_{i}$, along with a reformulation prompt $p_{2}$ (e.g., add essential scientific or mathematical explanation the search result would contain):
@@
Using only the subquestion $r_{i}$ to generate a query often results in noisy or overly specific formulations that include problem-specific variables or conditions irrelevant to general knowledge retrieval. On the other hand, using only the initial query $q_{i}$ lacks sufficient context to capture the intended reasoning path. By combining both, we enable the model to synthesize a logical query that is purpose-driven and aligned with the knowledge required to answer $r_{i}$. Even if the reformulated query $q_{i}^{*}$ contains factual inaccuracies, it tends to retrieve passages from a corpus $\mathcal{C}$ that are logically relevant and supportive of the reasoning required for solving the original problem. Figure 2 presents example queries generated by RAISE, Step-Back+RAG, and HyDE, illustrating RAISE’s ability to generate logical queries that are well-aligned with the reasoning intent.
Logical Retrieval.
The reformulated query $q_{i}^{*}$ is used to retrieve external knowledge:
@@
[…]
@@
GPQA SuperGPQA MMLU
Overall
science-
hard
science-
middle
engineering-
hard
Chemistry
Biology
College Chemistry
Direct
CoT 42.42 4.52 15.08 6.53 25.44 51.88 49.50
Direct+RAG
CoT+RAG 45.96 7.54 12.56 7.54 25.18 54.39 43.00
Decomposed
Least-to-Most 44.95 6.03 14.57 10.05 24.56 53.97 45.40
Step-Back 44.44 5.03 15.08 6.03 22.70 56.49 43.00
Decomposed+RAG
Least-to-Most+RAG 45.95 6.03 14.57 8.04 22.97 58.02 46.00
Step-Back+RAG 43.43 5.53 15.58 9.05 23.06 56.34 43.00
HyDE 46.46 7.54 13.07 7.04 22.97 57.88 49.00
Ours
RAISE
51.01
(+9.8%)
10.05
(+33.3%)
19.60
(+25.8%)
10.55
(+5.0%)
28.36
(+11.5%)
59.27
(+2.2%)
51.00
(+3.0%)
Table 1: Comparison of various reasoning strategies across GPQA, SuperGPQA, and MMLU. The underscore indicates the highest score among baselines, while the boldface indicates the highest score overall. The value in parentheses below the RAISE score represents the performance gain compared to the best baseline. RAISE consistently outperforms other approaches for scientific reasoning benchmarks.
@@ 4.1 Experimental Setup
4.1.1 Datasets.
To evaluate our method in settings that require multi-step scientific reasoning, we select three datasets focused on high-level STEM and professional reasoning. We provide more details on the datasets in Appendix B.
@@
  • SuperGPQA Team et al. (2025): This dataset evaluates the knowledge and reasoning abilities of LLMs across 285 graduate-level disciplines. We focus on three subsets that are challenging enough and require step-by-step reasoning in the science domain: science-hard, science-middle, and engineering-hard.
  • MMLU Hendrycks et al. (2021); Wang et al. (2024b): We focus on a subset of science-related tasks in MMLU-Pro, specifically chemistry and biology, which are well-suited for evaluating step-by-step reasoning with specialized knowledge. Given the lower overall difficulty of MMLU-STEM compared to MMLU-Pro, we additionally include its chemistry category as a lightweight validation setting.
4.1.2 Retriever Configuration.
We adopt Dense Passage Retrieval (DPR) Karpukhin et al. (2020) as our retriever, using the publicly available model trained on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
4.1.3 Language Model.
For GPQA, which is the primary focus of our study, we use Mistral Small 3.1-Instruct-2503 Mistral AI (2025), a 24B state-of-the-art small-scale open-sourced LLM that achieves strong performance on scientific reasoning tasks. For the remaining tasks, we use the smaller LLaMA 3.1-8B model Grattafiori et al. (2024) due to computational constraints, applying it to SuperGPQA, and MMLU. We provide detailed hyperparameters and experiment setups in Appendix C.
4.1.4 Baselines.
To assess the importance of multi-step reasoning and step-aware retrieval, we conduct experiments with four groups of baselines: Direct Reasoning, Direct Reasoning with RAG, Decomposed Reasoning, and Decomposed Reasoning with RAG.
Direct Reasoning solves problems in a single step without decomposing them into subquestions, and includes Chain-of-Thought (CoT) Wei et al. (2022), which generates intermediate reasoning steps through prompting. A retrieval-augmented variant, Direct Reasoning with RAG, includes CoT+RAG Lewis et al. (2020b), which retrieves documents using the original question as the query.
Decomposed Reasoning methods break down the original problem into subquestions and solve them sequentially. This includes Least-to-Most Zhou et al. (2023), which solves each subquestion directly, and Step-Back Zheng et al. (2024), which applies principle extraction at the subquestion level. A retrieval-augmented variant, Decomposed Reasoning with RAG, retrieves evidence for each subquestion and solves them step-by-step. This group includes Least-to-Most+RAG Lewis et al. (2020b), Step-Back+RAG Zheng et al. (2024), and HyDE Gao et al. (2023a), with the latter two enhancing retrieval relevance through query reformulation, making them strong baselines for comparison.
@@ 4.2 Main Results
4.2.1 Baseline Comparison.
As shown in Table 1, our proposed method, RAISE, consistently outperforms all baseline reasoning strategies across three benchmark datasets of varying difficulty: GPQA, SuperGPQA, and MMLU, achieving an average performance improvement of 13% over the best baseline scores. Unlike other baselines whose performance varies depending on the dataset’s difficulty or type, RAISE consistently demonstrates robust performance and outperforms them across different domains, types, and levels of difficulty. Specifically, on challenging GPQA, RAISE achieves a score of 51.01, outperforming the next-best method (HyDE) by 9.8%.
To assess the effectiveness of our logical query generation, we compare RAISE with three RAG-based decomposed reasoning baselines that differ in how they construct retrieval queries. Least-to-Most+RAG uses the subquestion itself as the query, Step-Back+RAG abstracts a general principle from the subquestion, and HyDE generates a hypothetical answer to use as the retrieval query. RAISE consistently outperforms all baselines across benchmarks, demonstrating the advantage of generating logically grounded queries that better align with the reasoning required to solve each subquestion. These results confirm that RAISE’s queries go beyond retrieving documents that are merely domain-relevant or superficially similar, enabling access to knowledge that is logically aligned with the problem-solving process.
@@
Unlike RAISE, decomposed reasoning methods do not always yield better performance, particularly for smaller LLMs that lack logically relevant knowledge Hosseini et al. (2024); Xu et al. (2025). While prior work Zhou et al. (2023) using decomposition focused on proprietary large models, our study targets open-source smaller LLMs. Without sufficient background knowledge, even well-structured subquestions can lead to hallucinations or unsupported reasoning, making decomposition potentially counterproductive unless relevant documents are retrieved.
@@
We also observe cases where models without retrieval outperform their retrieval-augmented variants. For example, on the SuperGPQA engineering-hard subset, Least-to-Most+RAG (8.04) performs worse than Least-to-Most (10.05). Although retrieval is intended to improve performance by providing additional information, naive retrieval can introduce noise, especially in open-domain settings where irrelevant content is common. Conventional RAG methods often retrieve documents that are only superficially related or worse, focus on irrelevant variables or distractors rather than the core principles required for reasoning. These results highlight the importance of retrieving logically relevant knowledge and further support the effectiveness of our approach.
4.2.2 Applying RAISE to various LLMs.
To assess the generalizability of RAISE across different LLM scales, we evaluate its performance on GPQA using LLaMA 3.1-8B Grattafiori et al. (2024) and GPT-4o mini OpenAI (2024), in addition to Mistral (used in our main experiments). As shown in Table 2, RAISE demonstrates consistent improvements over other baselines, exhibiting a similar trend to our main results with Mistral-24B. This shows that the effect of RAISE is not limited to a specific type of LLM, but can be applied to various LLMs with different scales.
@@
Step-Back 28.28 42.42 44.44
Decomposed+RAG
Least-to-Most+RAG 24.24 42.93 45.95
Step-Back+RAG 21.72 42.42 43.43
HyDE 25.75 38.89 46.46
Ours
RAISE
30.30
(+7.1%)
47.98
(+5.3%)
51.01
(+9.8%)
Table 2: Evaluation on GPQA with various LLMs with different scales: LLaMA 3.1-8B, GPT-4o mini, and Mistral Small 3.1. Underscore marks the best baseline; bold indicates the best overall. Values in parentheses under RAISE show gains over the top baseline. RAISE consistently shows the best performance across all settings.